# Language Models “Grok” to Copy

### v1

@@ Abstract
Ang Lv1,2, Ruobing Xie211footnotemark: 1,  Xingwu Sun2, Zhanhui Kang2, Rui Yan1thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China 2Machine Learning Platform Department, Tencent {anglv, ruiyan}@ruc.edu.cnruobingxie@tencent.com
@@ 1 Introduction
Large language models (LLMs) possess the capability to learn, retrieve, and reason from input context, facilitating various applications such as in-context learning (ICL, Brown et al., 2020) and retrieval-augmented generation (RAG, Lewis et al., 2020). Despite these achievements, several shortcomings have been reported regarding LLMs’ in-context capacities. For instance, the order of ICL demonstrations matters Lu et al. (2022); Zhang et al. (2024) and LLMs’ awareness of different contextual positions fluctuates Liu et al. (2023); Chen et al. (2024); Lin et al. (2024). We believe that studying the mechanisms behind the development of in-context capabilities during pre-training offers valuable insights for enhancing LLMs from a novel perspective.
@@
Our experimental method is summarized as follows: We trained 12-layer Llama models Touvron et al. (2023) using 40 billion tokens and saved checkpoints at regular intervals. To evaluate context copying, we presented the models with an input context comprising multiple random token subsequences, each beginning with a unique prefix, and let them to complete one of the prefixes presented in the context. The accuracy of these completions served as a measure of the models’ context copying abilities. By analyzing the evolution of context copying accuracy and the development of circuits (i.e., the subnetworks responsible for completing the specific task) across the saved checkpoints, we argue a potential connection between grokking and the development of context copying capabilities, as outlined in the following arguments:
@@
We observed that context copying accuracy shows a sudden increase long after the training loss stabilizes, akin to “grokking” on the test set when neural networks trained on small training sets.
@@
We noted that induction heads Olsson et al. (2022), attention heads responsible for copying tokens, form from shallow to deep layers during training, consistent with research showing deeper circuits form in Transformers after grokking Wang et al. (2024).
@@ 2 General Setup
We train small Llama models Touvron et al. (2023) on a subset of the RedPajama dataset Computer (2023), comprising 40 billion tokens, with the task of next-token prediction. Our model has 12 layers, each with 12 attention heads. The hidden state dimension is 768, and the intermediate dimension of MLP layers is 3,072. Our models contain 162M parameters. We use Llama tokenizer with a vocabulary of 32,000 tokens, and set the model context length to 1,024 tokens. Unless otherwise specified, the following hyperparameters are used: The AdamW optimizer Loshchilov and Hutter (2019) with $(\beta_{1},\beta_{2})$ = (0.9, 0.999), a learning rate of 0.1, 2000 warmup steps, and the norm clip value of 1. Our training is conducted on 8 A100 GPUs, with a batch size of 64 per GPU.
@@
The model’s objective is to continue the input. We consider the model’s output correct if it copies the suffix of the queried prefix from the context, since random token sequences lack meaningful semantics and the most natural continuation is to generate a prefix that has appeared in the context Olsson et al. (2022). To comprehensively assess context copying capabilities across different contextual positions, we evaluate the model for every $i$ mod $5=0$. We report the average accuracy across these positions. Unless specifically indicated, we report the average accuracy across these positions, from models trained with 3 different random seeds.
@@ 3 Language Models “Grok” to Copy
We propose that language models develop the context copying in a manner similar to “grokking”. In this section, we present three arguments and present the corresponding experiments and analyses that provide evidence for these claims.
@@
For argument 2, we trained another two models using the same setups and same initial weights as described in Section 2, but with batch sizes of 32 and 128. Our results indicate that grokked context copying is independent on the token count. Figure 3 shows that with a fixed learning rate, to achieve similar accuracy to models using a batch size of 64, models trained with a batch size of 128 (32) require twice (half) the token count, as their update steps are equal. This finding aligns with observations Wang et al. (2024) that data quantity does not affect the grokking speed. The consistency enhances the connection between grokking and the development of context copying.
@@
Moreover, we examine the impact of learning rates. Figure 5 indicates that an increased learning rate facilitates earlier and stronger grokking. Consequently, we assume that the grokked context copying emerges until the optimization reaches a specific intensity, which is influenced by both the learning rate and the number of update steps.
@@
Figure 6 illustrates the evolution of induction heads during training, revealing that these heads develop from shallower to deeper layers. This observation is consistent with Wang et al. (2024), which observes that after grokking, models develop circuits in deeper layers than before grokking.
@@ 4 Application
Viewing the development of context copying as a special grokking inspires us to examine the impact of regularization, as it is known to enhance grokking Wang et al. (2024); Nanda et al. (2023). We train models using (1) attention dropout with a 10% probability and (2) weight decay with $\lambda=0.1$. Figure 7 shows that these techniques’ positive impact: with dropout, the model groks to copy earlier; both techniques improve the accuracy compared to the vanilla model. Future innovations on improving the in-context learning could also benefit from the correlations with grokking.
@@ 5 Conclusions
We introduce a novel perspective that the development of context copying is a special grokking. It holds the potential to enhance the efficiency of language model training research because it enables us to explore grokking using small synthetic datasets, derive meaningful insights, and apply these insights to language models, as we did in Section 4. We hope a better understanding of grokking in future works provide more insights for developing language models with improved in-context capability.

### v2

@@ Abstract
Ang Lv1,2, Ruobing Xie211footnotemark: 1,  Xingwu Sun2, Zhanhui Kang2, Rui Yan1,3thanks:   Corresponding authors: Ruobing Xie (ruobingxie@tencent.com) and Rui Yan (ruiyan@ruc.edu.cn) 1Gaoling School of Artificial Intelligence, Renmin University of China 2Machine Learning Platform Department, Tencent 3School of Computer Science, Wuhan University {anglv, ruiyan}@ruc.edu.cnruobingxie@tencent.com
@@ 1 Introduction
Large language models (LLMs) can learn, retrieve, and reason from input context, facilitating various applications such as in-context learning (ICL, Brown et al., 2020) and retrieval-augmented generation (RAG, Lewis et al., 2020). Despite these achievements, several shortcomings have been reported regarding LLMs’ in-context capacities. For instance, the order of ICL demonstrations matters Lu et al. (2022) and LLMs’ awareness of different contextual positions fluctuates Liu et al. (2023). We believe that studying the mechanisms behind the development of in-context capabilities during pre-training offers valuable insights for enhancing LLMs from a novel perspective.
@@
Our experimental method is summarized as follows: We trained 12-layer Llama models Touvron et al. (2023) using 40 billion tokens and saved checkpoints at regular intervals. To evaluate context copying, we presented the models with an input context comprising multiple random token subsequences, each beginning with a unique prefix, and let them complete one of the prefixes presented in the context. The accuracy of these completions served as a measure of the models’ context copying abilities. By analyzing the evolution of context copying accuracy and the development of circuits (i.e., the subnetworks responsible for completing the specific task) across the saved checkpoints, we argue there is a potential connection between grokking and the development of context copying capabilities, as outlined in the following arguments:
@@
We observe that context copying accuracy shows a sudden increase long after the training loss stabilizes, akin to “grokking” on the test set when neural networks trained on small training sets.
@@
We note that induction heads Olsson et al. (2022), attention heads responsible for copying tokens, form from shallow to deep layers during training, consistent with research showing deeper circuits form in Transformers after grokking Wang et al. (2024).
@@ 2 General Setup
We train small Llama models Touvron et al. (2023) on a subset of the RedPajama dataset Computer (2023), comprising 40 billion tokens, with the task of next-token prediction. Our model has 162M parameters (12 layers, each with 12 attention heads; The hidden state dimension is 768, and the intermediate dimension of MLP layers is 3,072.) The context length is 1,024 tokens. We use the Llama tokenizer with a vocabulary of 32,000 tokens. Unless otherwise specified, the following hyperparameters are used: The AdamW optimizer Loshchilov and Hutter (2019) with $(\beta_{1},\beta_{2})$ = (0.9, 0.999), a learning rate of 0.1, 2000 warmup steps, and the norm clip value of 1. Our training is conducted on 8 A100 GPUs, with a batch size of 64 per GPU.
@@
We ask the model to continue the input. An output is correct if it copies the suffix of the queried prefix from the context, since random token sequences lack meaningful semantics and the most natural continuation is to generate the suffix of the prefix that has appeared in the context Olsson et al. (2022). To comprehensively assess context copying capabilities across different contextual positions, we evaluate the model for every $i$ mod $5=0$. Unless specifically indicated, we report the average accuracy across these positions, from models trained with 3 different random seeds.
@@ 3 Language Models “Grok” to Copy
We propose that language models develop context copying in a manner similar to “grokking”. This section presents three arguments, along with supporting experiments and analyses.
@@
For argument 2, we trained another two models using the same setups and same initial weights as described in Section 2, but with batch sizes of 32 and 128. Our results indicate that grokked context copying is independent of the token count. Figure 3 shows that with a fixed learning rate, to achieve similar accuracy to models using a batch size of 64, models trained with a batch size of 128 (32) require twice (half) the token count, as their update steps are equal. This finding aligns with observations Wang et al. (2024) that data quantity does not affect the grokking speed. The consistency enhances the connection between grokking and the development of context copying.
@@
Moreover, we examine the impact of learning rates. Figure 5 indicates that an increased learning rate facilitates earlier and stronger grokking. Consequently, we assume that the grokked context copying doesn’t emerge until the optimization reaches a specific intensity, which is influenced by both the learning rate and the number of update steps.
@@
Figure 6 illustrates the evolution of induction heads during training, revealing that they develop from shallower to deeper layers. This findings echos Wang et al. (2024), who proposes that after grokking, models develop circuits in deeper layers.
@@ 4 Application
Viewing the development of context copying as a special grokking inspires us to examine the impact of regularization, as it enhances grokking Nanda et al. (2023). We train models using (1) 10% attention dropout and (2) weight decay ($\lambda=0.1$). Figure 7 shows that their positive impact: with dropout, the model groks to copy earlier; both techniques improve the accuracy compared to the vanilla model.
5 Discussions
We sincerely appreciate the anonymous reviewers for their valuable feedback. In this section, we address key points raised in their reviews, which may also be of interest to a broader audience.
1. Our motivation for using copying tasks to measure in-context ability.
Induction heads, the key components responsible for in-context learning, are known to perform “copy and paste,” as described by Olsson et al. (2022). In essence, induction heads “complete the pattern” by copying and extending sequences that have occurred previously. This behavior motivates our exploration of copying, which are foundational to understanding in-context abilities.
Moreover, the copying task employed in this study has proven effective in previous research on RAG Tan et al. (2025) and in-context abilities Chen et al. (2024).
2. We suggest evaluating grokking through downstream performance rather than training loss.
In our task, the training objective is natural language modeling, while the testing task focuses on general copying. As a result, the training loss doesn’t fully capture the performance saturation seen in traditional grokking tasks. This is because copying can be viewed as a skill learned during pretraining, and once copying proficiency saturates, further improvements in other abilities can still lead to a decrease in training loss.
To demonstrate that copying on the training data has reached saturation, we measured the “ICL score” proposed by Olsson et al. (2022), which tracks the development of in-context abilities. Our results show that after approximately 4,000 training steps (about 1.85 billion tokens), the ICL score stabilizes at -0.5 nats. Since testing accuracy continues to improve well after this saturation point, we infer that once copying accuracy is “grokked,” reductions in training loss primarily stem from improvements in other abilities, rather than further progress in in-context copying.
3. The trade-offs between knowledge acquisition and in-context ability.
Some studies Chang et al. (2024) suggest that large batch sizes enhance knowledge acquisition but hinder the development of in-context abilities, highlighting a trade-off between the two Nafar et al. (2024); Yu et al. (2023). While large batch sizes slow down in-context ability acquisition, their overall effect in real-world applications remains difficult to quantify, necessitating further research.
4. Properties of Grokking
The properties of grokking are not limited to the three arguments we have exemplified. Many studies Miller et al. (2024); Fan et al. (2024); Liu et al. (2022); Lee et al. (2024) explore various aspects of grokking; we list some for readers who may be interested.
@@ 6 Conclusions
This paper introduces a novel perspective that the development of context copying is a special grokking. It holds the potential to provide meaningful insights that can be applied to language models, as we did in Section 4. We hope a better understanding of grokking in future works provide more insights for developing stronger language models.
Limitations
This paper focuses on the copying task to reflect the development of in-context capacities. Future innovations on improving the language model with better in-context capacities (e.g., ICL) might benefit from the correlations with grokking. However, it is important to note that ICL presents a higher level of complexity compared to simple copying tasks. Due to our limited computational resources, we were unable to train language models to achieve robust ICL performance, and therefore did not evaluate ICL tasks.
Acknowledgement
Ruobing Xie is supported by the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). Ang Lv is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2024 of Renmin University of China.